OpenStudy (anonymous):

Help? What is the denominator of the simplified form of 1 over 2 – 3 over 4 x minus 2 ? 2 4 x – 2 2x – 1

OpenStudy (anonymous):

I got A. but just incase

OpenStudy (mertsj):

\[\frac{1}{2}-\frac{3}{4x-2}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Look at it this way: \[\frac{ 1 }{ 2 }-\frac{ 3 }{ 2(2x-1) }\]

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@Luigi0210 , sorry i had to go to the store.